OpenStudy (anonymous):

An avg diameter of a star is aprox. 10^9 m and an average of a galaxy is 10^21 m. Considerng the given data, apply a sense of scale calculation to determine if when a galaxys diamete is scaled down to the lengeth of Manahattan (21.6 km) would the star be bigger or smaller than you

OpenStudy (anonymous):

so 10^9/x=10^21/21.6 x=2.16*10^-8 m that's quite smaller than you. :)